# Towards One-Stage End-to-End Table Structure Recognition with Parallel Regression for Diverse Scenarios

### v1

@@ Abstract
Keywords Table Structure Recognition One-stage End-to-End Parallel Regression
@@ 1 Introduction
Previously, most table structure recognition methods based on deep learning adopted a two-stage processing flow, such as [1, 2, 3] and other studies that first located the table region through object detection and subsequently conducted structural parsing on the cropped region, as illustrated in Figure 1(a). Although these methods can handle some standard document tables, they are difficult to cope with the problem of geometric deformation in natural scenes due to shooting angle, occlusion or bending. In recent years, some works have attempted to optimize the two-stage paradigm. For instance, TGRNet[4] and LORE[5] sequentially predicted the spatial location and logical indices of cells through two networks, as shown in Figure 1(b). While these methods perform well in different scenarios, their phased design requires independent training of multiple sub-modules to optimize performance, which significantly increases the training complexity. Furthermore, the serial inference process introduces additional computational overhead, making it difficult to meet the real-time requirements. Although the existing one-stage methods simplify the process, they face bottlenecks in scenario generalization. For example, TableNet[6] rely on the assumption of a clean and structured background and cannot handle blurring, uneven illumination, or geometric deformation in natural scenes. Although Cycle-CenterNet[7] and SCAN[8] support table parsing in natural scenes, it does not consider borderless tables and has insufficient accuracy in complex scenarios such as cross-row and cross-column merging. TRACE[9] is sensitive to geometric distortions such as bending and relies on complex post-processing algorithms. Overall, the existing methods have difficulty achieving an effective balance among cross-scenario adaptability, structural recognition robustness, and computational efficiency.
To address the above problems, we propose a one-stage end-to-end table structure parsing network, TableCenterNet, based on the CenterNet[10] framework. As illustrated in Figure 1(c), this method for the first time unifies the spatial location detection and logical location prediction of cells into a parallel regression task, and realizes efficient joint learning through the synergistic architecture of a shared feature extraction layer and task-specific decoding. This design not only avoids the error accumulation problem of the two-stage approach, but also significantly improves the parsing accuracy of complex tables (e.g., merging cells across rows and columns) by implicitly learning the spatial-logical location mapping law. Experiments show that TableCenterNet is highly competitive and outperforms existing state-of-the-art methods for logical location prediction on the TableGraph-24k[4] dataset.
@@ 2 Related Work
Early table structure recognition methods[11, 12, 13, 14, 15, 16] primarily rely on manually designed features and heuristic rules. For example, they parse tables using visual cues such as text arrangement, lines, and templates. Such methods are effective only for tables in specific formats (e.g., PDF) and struggle to handle complex tables. With the rise of statistical machine learning, approaches such as [17, 18, 19, 20] attempt to reduce the reliance on rules, but still require manual feature design, the stringent assumptions regarding table layouts restrict their practical applications. In recent years, deep learning-based methods have shown better performance than traditional methods[21]. Based on the descriptions of these methods, we can broadly categorize them into three groups: row/column division-based methods, markup sequence generation-based methods, and cell reconstruction-based methods.
2.1 Row/column division-based methods
These methods mainly reconstruct the table structure by detecting or segmenting row and column regions[1, 6, 22, 23, 24, 25, 26, 3], or by directly predicting the separation lines for rows and columns[27, 28, 29, 30, 31, 32, 9]. For example, DeepTabStR[22] improves the object detection model using deformable convolutions to directly locate rows and columns; DeepdeSRT[1] and TableNet[6] utilize semantic segmentation models to predict row and column regions and generate cells through intersection operations. However, none of the above methods can handle the cases of row-spanning and column-spanning. Therefore, SPLERGE[27] further proposes a two-stage strategy of segmentation and merging, using a segmentation model to generate an initial cell grid and then processing cross-cells through a merging module. But these methods rely on the strict alignment assumption of rows and columns and have poor adaptability to curved tables or non-axis-aligned layouts. Although RobusTabNet[32] enhances robustness to distorted tables by introducing a spatial convolution module, the heuristic rules in its post-processing stage may still falter due to inadequate segmentation quality.
@@ 2.2 Markup sequence generation-based methods
Such methods[33, 34, 35, 36, 37] treat table structure recognition as a generation task from images to tag sequences and directly output structured tags such as HTML or LaTeX. TableMaster[36] is built on the Transformer architecture, which extracts visual features through the encoder and generates HTML sequences via the decoder. TableFormer[37] not only decodes structural tags but also decodes cell coordinates. However, these methods usually rely on large-scale training data, and the length of the generated markup sequence increases sharply with greater table complexity, resulting in performance degradation. In addition, since the model uses a sequential decoding process, its inference is also more time-consuming.
@@ 2.3 Cell reconstruction-based methods
Cell reconstruction-based methods[38, 39, 40, 4, 41, 7, 42, 43, 44, 8, 5] usually predicts table cells first and then utilize these cells as the basic units to construct the global table structure by analyzing the associative information between them. For example, GraphTSR[38] models the predicted cells as graph nodes and employs a graph attention network to predict horizontal, vertical, or irrelevant relationships. FLAG-Net[40] enhances the reasoning ability for cell associations through both dense and sparse context modeling. TGRNet[4] and LORE[5] perform logical location prediction on the detected cells using Graph Convolutional Networks and Cascade Transformers, respectively, and directly generate logical indices. Existing methods generally adopt a two-stage process of "detection-association". Our method, however, makes full use of the cell location distribution information in the feature map during the detection stage, performs parallel regression of spatial and logical location, and directly matches the physical coordinates with the logical location through interpolation maps to obtain the logical indices, thus simplifying the two-stage process into a one-stage end-to-end parsing.
@@ 3.1 Overall Architecture
The overview of the proposed one-stage TableCenterNet is shown in Figure 2. It employs a CNN backbone to extract visual features of cells from the input image, and then jointly predicts the spatial locations and logical locations of cells in a multi-task manner through six parallel regression heads. is the number of cells, and the definitions of and are as follows:
Where denotes the coordinates of the -th corner point of the -th cell. The four corner points are arranged clockwise starting from the upper-left corner. , , , and correspond to the four logical indices of the -th cell: start row, end row, start column, and end column, respectively.
@@ 3.2 Spatial Location Regression
To enhance the accuracy of spatial location prediction, we adopted the Cycle-Pairing Module[7]. By regressing the vectors from the center of each cell to the four corner points, as well as the vectors from the four corner points back to the center, and combining the corresponding cell center coordinates , we can calculate the approximate spatial locations of the cells. The specific calculation formula is as follows:
In addition, we also improved the PairLoss[7] to better combine it with logical location regression and to minimize the interference caused by invalid vectors from the corners to the center. The formula for the new spatial location regression loss is as follows:
Among them, and represent the GT corresponding to the predicted vectors and , respectively. denotes the MAE loss function, while represents the loss of the vector from the invalid corner point to the center. The hyperparameters , , and are used to adjust the relative importance of these loss items. Additionally, dynamically weights the loss based on the quality of the regression, it is defined as follows:
Taking corner points as reference points, each corner point may point to the centers of at most four different cells, i.e., corner-to-center vectors can be obtained through regression, where is the total number of corner points. Since , in addition to the vectors matched by cells, there are also many invalid corner-to-center vectors, which are defined as . The loss is formulated as follows:
@@ 3.3 Logical Location Regression
Considering that the extracted feature maps already contain the location information of cells, we capture the row-column demarcation features of the table by regressing the corresponding logical location interpolation map, obtaining the row interpolation map and the column interpolation map . Meanwhile, based on the cell center, we regress the row-column span of each cell, with the purpose of synergistically supervise the regression of the interpolation map and simplify the parsing process.
@@
Then, we input and into Algorithm 1 to calculate the corresponding row interpolation map and column interpolation map . Since the physical coordinates of the input cells are the same in the two calculations, the obtained mask image is also the same. The visualization results are shown in Figure 3. During training, we only need to reduce , , and by a factor of 4 to use them as the GT for interpolation map regression.
@@
Where is a mapping function used to enhance the supervision of cell boundaries, which can be defined as follows:
Supervision of cell spans stipulates that the logical location of each cell should be consistent with its span. Formally, the loss for constraining cell spans can be defined as follows:
@@ 3.4 Parsing Processing
In Figure 4(a), obtained through regression are only approximations. The detected cell corner points (e.g., Figure 4(b)) need to be used for alignment processing[7] to obtain more accurate spatial locations , as shown in Figure 4(c). Then, calculate the physical coordinates of all cells through 4x magnification.
Subsequently, by directly matching through physical coordinates and the interpolation map of logical location, the logical indices of all cells can be obtained. The specific matching process for each cell is shown in the following formula:
@@ 3.5 Training Strategy
Selective gridding interpolation. For datasets with unbiased prior knowledge of physical and logical structures and no significant deformation, there is no need to generate interpolation maps based on cells before training. As shown in Figure 5, the cells can be converted into a logical grid structure, i.e., decomposing the cross-row and cross-column cells into multiple logical grid cells. This effectively unifies the interpolation style of each cell and improves the regression accuracy of the interpolation map.
@@ 4.3 Implementation Details
All experiments were conducted in a PyTorch environment and executed on a workstation equipped with two GeForce RTX 3090 GPUs. In these experiments, we respectively adopted DLA-34[50] and StartNet-s3[51] as the backbone networks, and their weights were all initialized based on pre-trained models for the ImageNet[52] classification task. Within the CenterNet[10] framework’s regression network, we configured the number of hidden channels for each convolutional head to . Model optimization was performed using the Adam algorithm with a batch size of . For both backbone networks, the models were trained for epochs, with the initial learning rate set to . The learning rate was decayed to and in the 140th and 180th epochs, respectively. To ensure scale invariance, we uniformly resized the input table images to a fixed size of pixels, and pixels for the TableGraph-24k dataset, while maintaining the aspect ratio of the images. After resizing, we also applied data augmentation and normalization to the input images to accelerate the convergence of the network.
Due to the limited size of the ICDAR-2013 dataset, which contains only 258 table images after preprocessing, we opted to perform rounds of fine-tuning on the model trained on WTW using the segmented ICDAR-2013 training set. During the fine-tuning phase, we expanded the original image by adding 100 black pixels to each of the four directions: top, bottom, left, and right, before scaling the image. The initial learning rate decays to and in the 70th and 90th epochs, respectively, while ensuring that the other training configurations remain unchanged. In the testing phase, we similarly expand the perimeter of the test image by black pixels before performing inference.
@@ 4.4 Results
To evaluate TableCenterNet’s ability to recognize table structures in different scenarios, we tested it on three benchmark datasets with different scenarios. The qualitative results in Figure 6 show that our method is robust to complex structures, diverse colors, geometric transformations, and distorted or deformed tables in both photographic and scanning scenes. Furthermore, our method also demonstrates generalization to digital native scenarios, effectively handling tables across cells, with bounded cells, and without bounded cells.
@@
| ReS2TIM[48] | - | ICDAR2013† | - | - | - | 73.4 | 74.7 | 74.0 | 17.4 |
| TGRNet[4] | ResNet-50 | ICDAR2013† | 68.2 | 65.2 | 66.7 | - | - | - | 27.5 |
@@
| LGPMA[25] | ResNet-50 | ICDAR2013† | - | - | - | 96.7 | 99.1 | 97.9 | - |
| Cycle-CenterNet*[7] | DLA-34 | WTW | - | - | - | 97.5 | 98.4 | 98.0 | - |
@@
| SCAN*[8] | ResNet-50 | WTW | - | - | - | 98.2 | 98.1 | 98.2 | - |
| FLAGNet[40] | ResNet-50 | ICDAR2013† | - | - | - | 97.9 | 99.3 | 98.6 | - |
| NCGM[42] | ResNet-18 | ICDAR2013† | - | - | - | 98.4 | 99.3 | 98.8 | - |
| LORE[5] | DLA-34 | ICDAR2013† | - | - | 97.2 | 99.2 | 98.6 | 98.9 | 86.8 |
| TableCenterNet-D | DLA-34 | ICDAR2013† | 98.7 | 96.9 | 97.8 | 96.1 | 99.1 | 97.6 | 91.1 |
| TableCenterNet-D* | DLA-34 | WTW | 97.1 | 97.9 | 97.5 | 98.4 | 98.6 | 98.5 | 94.6 |
| TableCenterNet-S | StartNet-s3 | ICDAR2013† | 98.2 | 96.8 | 97.5 | 93.7 | 97.5 | 95.6 | 88.1 |
| TableCenterNet-S* | StartNet-s3 | WTW | 97.1 | 98.2 | 97.6 | 98.4 | 98.6 | 98.5 | 94.1 |
@@ 4.5 Ablation Study
Effectiveness of cell boundaries and spans supervisions. Based on the results of Exp.1a and Exp.1b, we observe that the cell-span regression yields a performance improvement of +4.4% Acc. Comparing Exp.1a and Exp.1c, the cell span supervised loss leads to an even more substantial enhancement, with a performance improvement of +6.5% Acc. This is due to the fact that provides anchors for interpolation plot regression and strengthens the constraints on the range of interpolation between cells. From Exp.1d and Exp.1f, the cell boundary supervised loss also improves by +0.6%Acc due to the fact that it reduces the interpolation interference. According to the properties of interpolation map generation, the logical indices of cell boundaries are closer to the true values, while the logical indices of in-cell interpolation are approximate and less reliable. In Figure 7, we visualize the accuracy of each logical location by heat map. From the three heat maps, we can find that cell spanning regression improves the accuracy of small logical indices. When the two supervision mechanisms are combined, the accuracy of large logical indices is also significantly improved.
Effect of different logical location regression methods. In Exp.2a and Exp.2b, we replace interpolation map regression with upper-left corner point regression and four corner points regression, respectively, and both of them use the optimal loss computation scheme, and the network structure remains consistent with TableCenterNet. Compared with Exp.2f, the performance of these two logical location regression methods decreases by 3.2%Acc and 1.2%Acc, respectively, indicating that the interpolation map regression can better model the dependencies and constraints between logical locations of different cells. When there is an offset in the spatial locations regressed by the methods in Exp. 2a and Exp. 2b, the logical locations are prone to errors, whereas interpolation maps can ensure the accuracy of logical indices as much as possible by rounding the interpolation results.
| Exp. | Regression | Objectives | Spatial Location | Logical Location | ||||||
| Logical Location | Cell Span | P | R | F1 | A-c | A-r | Acc | |||

### v2

# TableCenterNet: A one-stage network for table structure recognition

@@ Abstract
Keywords Document Understanding Table Structure Recognition One-stage End-to-End
@@ 1 Introduction
Previously, most table structure recognition (TSR) methods based on deep learning adopted a two-stage processing flow. For example, studies such as [1, 2, 3] first located the table region through object detection and then performed structural parsing on the cropped region, as illustrated in 1(a). Although these methods can handle some standard document tables, they are difficult to cope with the problem of geometric deformation in natural scenes due to shooting angle, occlusion or bending. In recent years, some works have attempted to optimize the two-stage paradigm. For instance, TGRNet [4] and LORE [5] sequentially predicted the spatial location and logical indices of cells through two networks, as shown in 1(b). While these methods perform well in different scenarios, their phased design requires independent training of multiple sub-modules to optimize performance, which significantly increases the training complexity. Furthermore, the serial inference process introduces additional computational overhead, making it difficult to meet the real-time requirements. Although the existing one-stage methods simplify the process, they face bottlenecks in scenario generalization. For example, TableNet [6] rely on the assumption of a clean and structured background and cannot handle blurring, uneven illumination, or geometric deformation in natural scenes. Although Cycle-CenterNet [7] and SCAN [8] support table parsing in natural scenes, it does not consider borderless tables and has insufficient accuracy in complex scenarios such as cross-row and cross-column merging. TRACE [9] is sensitive to geometric distortions such as bending and relies on complex post-processing algorithms. Overall, existing methods struggle to achieve an effective balance between cross-scenario adaptability, structural recognition robustness, and computational efficiency.
To address the above problems, we propose a one-stage end-to-end table structure parsing network, TableCenterNet, based on the CenterNet [10] framework. As illustrated in 1(c), this method unifies the spatial location detection and logical location prediction of cells into a parallel regression task for the first time, and realizes efficient joint learning through the synergistic architecture of a shared feature extraction layer and task-specific decoding. This design not only avoids the error accumulation problem of the two-stage approach, but also significantly improves the parsing accuracy of complex tables (e.g., merging cells across rows and columns) by implicitly learning the spatial-logical location mapping law. Experiments show that TableCenterNet is highly competitive and outperforms existing state-of-the-art methods for logical location prediction on the TableGraph-24k [4] dataset.
@@ 2 Related Work
Early TSR methods [11, 12, 13, 14, 15, 16] primarily rely on manually designed features and heuristic rules. For example, they parse tables using visual cues such as text arrangement, lines, and templates. These methods are effective only for tables in specific formats (e.g., PDF) and struggle to handle complex tables. With the rise of statistical machine learning, approaches such as [17, 18, 19, 20] attempt to reduce the reliance on rules, but still require manual feature design, the strong assumptions regarding table layouts restrict their practical applications. In recent years, deep learning-based methods have shown better performance than traditional methods [21]. Based on the descriptions of these methods, we can broadly categorize them into three groups: row and column division-based methods, markup sequence generation-based methods, and cell reconstruction-based methods.
2.1 Row and column division-based methods
These methods mainly reconstruct the table structure by detecting or segmenting row and column regions [1, 6, 22, 23, 24, 25, 26, 3], or by directly predicting the separation lines for rows and columns [27, 28, 29, 30, 31, 32, 9]. For example, DeepTabStR [22] improves the object detection model using deformable convolutions to directly locate rows and columns; DeepdeSRT [1] and TableNet [6] utilize semantic segmentation models to predict row and column regions and generate cells through intersection operations. However, none of the above methods can handle the cases of row-spanning and column-spanning. Therefore, SPLERGE [27] further proposes a two-stage strategy of segmentation and merging, using a segmentation model to generate an initial cell grid and then processing cross-cells through a merging module. But these methods rely on the strict alignment assumption of rows and columns and have poor adaptability to curved tables or non-axis-aligned layouts. Although RobusTabNet [32] enhances robustness to distorted tables by introducing a spatial CNN module [33], the heuristic rules in its post-processing stage may still affect the overall performance due to low-quality separator masks.
@@ 2.2 Markup sequence generation-based methods
Such methods [34, 35, 36, 37, 38] treat TSR as a generation task from images to tag sequences and directly output structured tags such as HTML or LaTeX. TableMaster [37] is built on the Transformer architecture, which extracts visual features through the encoder and generates HTML sequences via the decoder. TableFormer [38] not only decodes structural tags but also decodes cell coordinates. However, these methods usually rely on large-scale training data, and the length of the generated markup sequence increases with greater table complexity, resulting in performance degradation. In addition, since the model uses a sequential decoding process, its inference is also more time-consuming.
@@ 2.3 Cell reconstruction-based methods
Cell reconstruction-based methods [39, 40, 41, 4, 42, 7, 43, 44, 45, 8, 5] usually predicts table cells first and then utilize these cells as the basic units to construct the global table structure by analyzing the associative information between them. For example, GraphTSR [39] models the predicted cells as graph nodes and employs a graph attention network to predict horizontal, vertical, or irrelevant relationships. FLAG-Net [41] enhances the reasoning ability for cell associations through both dense and sparse context modeling. TGRNet [4] and LORE [5] perform logical location prediction on the detected cells using Graph Convolutional Networks and Cascade Transformers, respectively, and directly generate logical indices. Existing methods generally adopt a two-stage process of "detection-association". Our method, however, makes full use of the cell location distribution information in the feature map during the detection stage, performs parallel regression of spatial and logical location, and directly matches the physical coordinates with the logical locations through interpolation maps to obtain the logical indices, thus simplifying the two-stage process into a one-stage end-to-end parsing.
@@ 3.1 Overall Architecture
The overview of the proposed one-stage TableCenterNet is shown in Fig. 2. It employs a CNN backbone to extract visual features of cells from the input image, and then jointly predicts the spatial locations and logical locations of cells in a multi-task manner through six parallel regression heads. is the number of cells, and the definitions of and are as follows:
Where denotes the coordinates of the -th corner point of the -th cell. The four corner points are arranged clockwise starting from the upper-left corner. , , , and correspond to the four logical indices of the -th cell: starting-row, ending-row, starting-column, and ending-column, respectively.
@@ 3.2 Spatial Location Regression
To enhance the accuracy of spatial location prediction, we adopted the Cycle-Pairing Module [7]. By regressing the vectors from the center of each cell to its four corners (center-to-corners) and the vectors from the four corners to the center (corners-to-center), and based on the center coordinates of each cell, we can calculate the approximate spatial locations of the cells. The specific calculation formula is as follows:
In addition, we have optimized PairLoss [7] to work more effectively in conjunction with logical position regression, while reducing the interference caused by invalid corners-to-center vectors. The formula for the improved spatial location regression loss is as follows:
Among them, and represent the ground truth (GT) corresponding to the predicted vectors and , respectively. denotes the mean absolute error [46] loss function, while represents the loss of invalid corners-to-center vectors. The hyperparameters , , and are used to adjust the relative importance of these loss items. Additionally, dynamically weights the loss based on the quality of the regression, it is defined as follows:
Using the corners as reference points, each corner can point to the centers of up to four different cells, resulting in corners-to-center vectors obtained through regression, where denotes the total number of cell corners. Since , in addition to the vectors matched by cells, there are also many invalid corners-to-center vectors, which are defined as . From this, the loss is formulated as follows:
@@ 3.3 Logical Location Regression
Considering that the extracted feature maps already contain the location information of cells, we capture the row-column demarcation features of the table by regressing the corresponding logical location interpolation map, obtaining the row interpolation map and the column interpolation map . Meanwhile, based on the cell center, we regress the row-column span of each cell, with the purpose of synergistically supervise the regression of the interpolation map and simplify the parsing process. Among them, and denote the row and column span of the -th cell respectively.
@@
Then, we input and into Alg. 1 to calculate the corresponding row interpolation map and column interpolation map . Since the physical coordinates of the input cells are the same in the two calculations, the obtained mask image is also the same. The visualization results are shown in Fig. 3. During training, we only need to reduce , , and by a factor of 4 to use them as the GT for interpolation map regression.
@@
Where and are weight functions used to enhance the supervision of the row-column demarcation of the table. In the interpolation map, the closer to the row-column demarcation, the higher the confidence of logical indices, and thus the weight should increase accordingly. and can be defined as follows:
Supervision of cell spans stipulates that the logical indices of each cell in the interpolation map should be consistent with its spans. Formally, the loss for constraining cell spans can be defined as follows:
@@ 3.4 Parsing Processing
In 4(a), obtained through regression are only approximations. The detected cell corner points (see 4(b)) need to be used for alignment processing [7] to obtain more accurate spatial locations , as shown in 4(c). Then, calculate the physical coordinates of all cells through 4x magnification.
Subsequently, by directly matching the physical coordinates with the interpolation maps, the logical indices of all cells can be obtained. The matching process for the logical indices of each cell can be formulated as follows:
@@ 3.5 Training Strategy
Selective gridding interpolation. For datasets with unbiased prior knowledge of physical and logical structures and no significant deformation, there is no need to generate interpolation maps based on cells before training. As shown in Fig. 5, the cells can be converted into a logical grid structure, i.e., decomposing the cross-row and cross-column cells into multiple logical grid cells. This effectively unifies the interpolation style of each cell and improves the regression accuracy of the interpolation map. Instead of using cells to generate interpolation maps before training, grids are used.
@@ 4.3 Implementation Details
All experiments were conducted in a PyTorch environment and executed on a workstation equipped with two GeForce RTX 3090 GPUs. In these experiments, we respectively adopted DLA-34 [52] and StartNet-s3 [53] as the backbone networks, and their weights were all initialized based on pre-trained models for the ImageNet [54] classification task. Within the CenterNet [10] framework’s regression network, we configured the number of hidden channels for each regression head to . Model optimization was performed using the Adam algorithm with a batch size of ( for TableGraph-24k). For both backbone networks, the models were trained for epochs, with the initial learning rate set to . The learning rate was decayed to and in the 140th and 180th epochs, respectively. To ensure scale invariance, we uniformly resized the input table images to a fixed size of pixels, and pixels for the TableGraph-24k dataset, while maintaining the aspect ratio of the images. After resizing, we also applied data augmentation and normalization to the input images to accelerate the convergence of the network.
Due to the limited size of the ICDAR-2013 dataset, which contains only 258 table images after preprocessing, we opted to perform rounds of fine-tuning on the model trained on WTW using the segmented ICDAR-2013 training set. During the fine-tuning phase, considering that most tables in the WTW dataset have outer margins, we extend the original image by 100 black pixels on all sides before scaling the image. The initial learning rate decays to and in the 70th and 90th epochs, respectively, while ensuring that the other training configurations remain unchanged. In the testing phase, we similarly expand the perimeter of the test image by black pixels before performing inference.
@@ 4.4 Results
To evaluate TableCenterNet’s ability to recognize table structures in different scenarios, we tested it on three benchmark datasets with different scenarios. The qualitative results in Fig. 6 show that our method is robust to complex structures, diverse colors, geometric transformations, and distorted or deformed tables in both photographic and scanning scenes. Furthermore, our method also demonstrates generalization to digital native scenarios, effectively handling tables across cells, with bounded cells, and without bounded cells.
@@
ReS2TIM [50] | - | ICDAR2013 | - | - | - | 73.4 | 74.7 | 74.0 | 17.4 |
TGRNet [4] | ResNet-50 | ICDAR2013 | 68.2 | 65.2 | 66.7 | - | - | - | 27.5 |
@@
LGPMA [25] | ResNet-50 | ICDAR2013 | - | - | - | 96.7 | 99.1 | 97.9 | - |
Cycle-CenterNet [7] | DLA-34 | WTW | - | - | - | 97.5 | 98.4 | 98.0 | - |
@@
SCAN [8] | ResNet-50 | WTW | - | - | - | 98.2 | 98.1 | 98.2 | - |
FLAGNet [41] | ResNet-50 | ICDAR2013 | - | - | - | 97.9 | 99.3 | 98.6 | - |
NCGM [43] | ResNet-18 | ICDAR2013 | - | - | - | 98.4 | 99.3 | 98.8 | - |
LORE [5] | DLA-34 | ICDAR2013 | - | - | 97.2 | 99.2 | 98.6 | 98.9 | 86.8 |
TableCenterNet-D | DLA-34 | ICDAR2013 | 98.7 | 96.9 | 97.8 | 96.1 | 99.1 | 97.6 | 91.1 |
TableCenterNet-D | DLA-34 | WTW | 97.1 | 97.9 | 97.5 | 98.4 | 98.6 | 98.5 | 94.6 |
TableCenterNet-S | StartNet-s3 | ICDAR2013 | 98.2 | 96.8 | 97.5 | 93.7 | 97.5 | 95.6 | 88.1 |
TableCenterNet-S | StartNet-s3 | WTW | 97.1 | 98.2 | 97.6 | 98.4 | 98.6 | 98.5 | 94.1 |
@@ 4.5 Ablation Study
Effectiveness of cell boundaries and spans supervisions. Based on the results of Exp.1a and Exp.1b, we observe that the cell-span regression yields a performance improvement of +4.4% Acc. Comparing Exp.1a and Exp.1c, the cell span supervised loss leads to an even more substantial enhancement, with a performance improvement of +6.5% Acc. This is due to the fact that provides anchors for interpolation plot regression and strengthens the constraints on the range of interpolation between cells. From Exp.1d and Exp.1f, the cell boundary supervised loss also improves by +0.6%Acc due to the fact that it reduces the interpolation interference. According to the properties of interpolation map generation, the logical indices of cell boundaries are closer to the true values, while the logical indices of in-cell interpolation are approximate and less reliable. In Fig. 7, we visualize the accuracy of each logical location by heat map. From the three heat maps, we can find that cell spanning regression improves the accuracy of small logical indices. When the two supervision mechanisms are combined, the accuracy of large logical indices is also significantly improved.
Impact of different logical location regression methods. In Exp.2a and Exp.2b, we replace interpolation map regression with logical indices regression of the upper-left corner point and the four corner points respectively. Both of them use the optimal loss computation scheme, and the network structure remains consistent with TableCenterNet. Compared with Exp.2f, the performance of these two logical location regression methods decreases by 3.2%Acc and 1.2%Acc, respectively, indicating that the interpolation map regression can better model the dependencies and constraints between logical locations of different cells. When there is an offset in the spatial locations regressed by the methods in Exp.2a and Exp.2b, the logical locations are prone to errors, whereas interpolation maps can ensure the accuracy of logical indices as much as possible by rounding the interpolation results.
Exp. | Regression | Supervisions | Spatial Location | Logical Location | ||||||
Logical Indices | Cell Span | P | R | F1 | A-c | A-r | Acc | |||